OpenStudy (tanabugg):

Find the surface area and volume of a cylinder with a radius of 4 and a height of 6 (pls explain the answer)

jhonyy9 (jhonyy9):

do you know formula of volume and surface area ?

OpenStudy (jdoe0001):

\(\textit{volume of a cylinder}=\pi rh\qquad \qquad \begin{cases} r=radius\\ h=height \end{cases} \\ \quad \\ \textit{surface area of a cylinder}=2\pi r(h+r)\)

jhonyy9 (jhonyy9):

surface area = circumference time height volume = area of base time height

jhonyy9 (jhonyy9):

@jdoe0001 you have missed there anything - check it please

OpenStudy (tanabugg):

The area of one circle would be 50.24 right?

jhonyy9 (jhonyy9):

than 50,24=3,14*4^2 so than yes

OpenStudy (tanabugg):

And both bases would be 100.48?